# Multivariate Variational Autoencoder

### v1

@@ Abstract
We present the Multivariate Variational Autoencoder (MVAE), a VAE variant that preserves Gaussian tractability while lifting the diagonal posterior restriction. MVAE factorizes each posterior covariance, where a global coupling matrix $\mathbf{C}$ induces dataset-wide latent correlations and per-sample diagonal scales modulate local uncertainty. This yields a full-covariance family with analytic KL and an efficient reparameterization via $\mathbf{L}=\mathbf{C}\mathrm{diag}(\boldsymbol{\sigma})$. Across Larochelle-style MNIST variants, Fashion-MNIST, CIFAR-10, and CIFAR-100, MVAE consistently matches or improves reconstruction (MSE $\downarrow$) and delivers robust gains in calibration (NLL/Brier/ECE $\downarrow$) and unsupervised structure (NMI/ARI $\uparrow$) relative to diagonal-covariance VAEs with matched capacity, especially at mid-range latent sizes. Latent-plane visualizations further indicate smoother, more coherent factor traversals and sharper local detail. We release a fully reproducible implementation with training/evaluation scripts and sweep utilities to facilitate fair comparison and reuse.
@@ V-A Experimental Setup
Each configuration was defined by a dataset–model–latent-size triplet, ensuring a systematic exploration of both the baseline and proposed architectures. For every dataset, we trained models with latent dimensions ranging from low to high capacity (typically $d_{z}\in\{2,4,16,32,256,512\}$), enabling a consistent comparison across scales of representation.
Training was performed for $200$ epochs using the AdamW optimizer with a learning rate of $10^{-3}$ and a batch size of $100$. All experiments used fixed random seeds and stratified splits to ensure reproducibility. Evaluation was based on a comprehensive set of quantitative metrics: reconstruction quality via per-pixel mean squared error (MSE), generative quality via the evidence lower bound (ELBO), latent linear separability using a logistic regression probe, calibration performance through negative log-likelihood (NLL), Brier score, and expected calibration error (ECE), as well as unsupervised clustering measures using normalized mutual information (NMI) and adjusted Rand index (ARI).
Each model is trained for $200$ epochs using the AdamW optimizer with a learning rate of $10^{-3}$ and a batch size of $100$. Datasets include the Larochelle MNIST variants (basic, rotated, background images, background random), Fashion-MNIST, CIFAR-10, and CIFAR-100. For each dataset, both VAE and MVAE models are trained with latent sizes $\{2,4,16,32,256,512\}$ depending on input dimensionality.
@@
Latent MSE/test $\downarrow$ Probe Acc. $\uparrow$ Probe NLL $\downarrow$ Brier $\downarrow$ ECE $\downarrow$ NMI $\uparrow$ ARI $\uparrow$
VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE
2 0.036 0.035 0.221 0.213 2.174 1.889 0.871 0.870 0.046 0.036 0.073 0.079 0.034 0.037
4 0.029 0.029 0.265 0.259 2.065 2.067 0.842 0.843 0.037 0.030 0.093 0.091 0.050 0.047
16 0.019 0.020 0.326 0.341 1.899 1.889 0.797 0.793 0.030 0.038 0.097 0.106 0.055 0.058
32 0.015 0.015 0.373 0.370 1.800 1.810 0.766 0.769 0.030 0.034 0.109 0.113 0.063 0.061
256 0.015 0.015 0.425 0.419 1.639 1.659 0.714 0.720 0.033 0.030 0.106 0.111 0.057 0.059
512 0.015 0.016 0.430 0.420 1.628 1.679 0.709 0.725 0.026 0.031 0.119 0.117 0.063 0.062
@@
TABLE I: VAE vs MVAE on Cifar10 across latent sizes. Best per cell is bold.
Latent MSE/test $\downarrow$ Probe Acc. $\uparrow$ Probe NLL $\downarrow$ Brier $\downarrow$ ECE $\downarrow$ NMI $\uparrow$ ARI $\uparrow$
VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE
2 0.037 0.037 0.037 0.036 4.423 4.429 0.984 0.985 0.008 0.007 0.174 0.175 0.012 0.012
4 0.029 0.030 0.075 0.077 4.162 4.147 0.973 0.973 0.019 0.020 0.203 0.206 0.018 0.019
16 0.018 0.019 0.137 0.139 3.821 3.821 0.945 0.946 0.020 0.024 0.219 0.217 0.026 0.026
32 0.015 0.016 0.165 0.169 3.667 3.660 0.930 0.930 0.017 0.021 0.225 0.222 0.028 0.027
256 0.016 0.016 0.199 0.202 3.471 3.469 0.908 0.907 0.025 0.023 0.223 0.224 0.027 0.028
512 0.015 0.016 0.198 0.194 3.509 3.505 0.913 0.912 0.031 0.026 0.223 0.225 0.027 0.026
@@
TABLE II: VAE vs MVAE on Cifar100 across latent sizes. Best per cell is bold.
Latent MSE/test $\downarrow$ Probe Acc. $\uparrow$ Probe NLL $\downarrow$ Brier $\downarrow$ ECE $\downarrow$ NMI $\uparrow$ ARI $\uparrow$
VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE
2 0.014 0.015 0.804 0.797 0.542 0.548 0.275 0.280 0.022 0.018 0.558 0.535 0.425 0.379
4 0.015 0.015 0.863 0.865 0.399 0.385 0.201 0.195 0.034 0.026 0.502 0.528 0.365 0.374
16 0.027 0.027 0.640 0.670 1.097 0.959 0.498 0.447 0.102 0.069 0.531 0.544 0.356 0.374
32 0.015 0.014 0.819 0.815 0.510 0.506 0.259 0.257 0.027 0.025 0.527 0.523 0.384 0.377
256 0.019 0.019 0.754 0.724 0.701 0.767 0.350 0.381 0.030 0.024 0.589 0.560 0.450 0.400
512 0.016 0.016 0.863 0.865 0.389 0.381 0.197 0.194 0.026 0.026 0.526 0.541 0.353 0.369
@@
TABLE III: VAE vs MVAE on Fashion MNIST across latent sizes. Best per cell is bold.
Latent MSE/test $\downarrow$ Probe Acc. $\uparrow$ Probe NLL $\downarrow$ Brier $\downarrow$ ECE $\downarrow$ NMI $\uparrow$ ARI $\uparrow$
VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE
2 0.0121 0.0119 0.904 0.910 0.308 0.298 0.141 0.136 0.0147 0.0137 0.623 0.641 0.557 0.579
4 0.0119 0.0120 0.943 0.951 0.192 0.168 0.087 0.076 0.021 0.019 0.598 0.623 0.527 0.552
16 0.0394 0.0398 0.605 0.561 1.145 1.215 0.516 0.554 0.083 0.076 0.525 0.514 0.389 0.376
32 0.0114 0.0111 0.913 0.924 0.287 0.254 0.130 0.115 0.018 0.013 0.589 0.642 0.497 0.582
256 0.0280 0.0277 0.860 0.864 0.572 0.562 0.232 0.227 0.073 0.074 0.571 0.582 0.443 0.456
512 0.0126 0.0128 0.945 0.951 0.187 0.175 0.084 0.078 0.023 0.022 0.572 0.602 0.481 0.528
TABLE IV: VAE vs MVAE on MNIST Basic across runs. Best per cell is bold.
Latent MSE/test $\downarrow$ Probe Acc. $\uparrow$ Probe NLL $\downarrow$ Brier $\downarrow$ ECE $\downarrow$ NMI $\uparrow$ ARI $\uparrow$
VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE
2 0.016 0.016 0.486 0.494 1.391 1.387 0.632 0.628 0.021 0.021 0.135 0.143 0.071 0.074
4 0.014 0.014 0.677 0.701 0.989 0.923 0.450 0.421 0.080 0.073 0.125 0.098 0.065 0.052
16 0.049 0.049 0.246 0.197 2.011 2.190 0.804 0.862 0.032 0.050 0.159 0.165 0.061 0.119
32 0.013 0.013 0.525 0.531 1.320 1.305 0.597 0.590 0.022 0.022 0.112 0.122 0.056 0.063
256 0.038 0.038 0.334 0.336 1.693 1.691 0.730 0.730 0.022 0.017 0.175 0.204 0.097 0.111
512 0.014 0.014 0.697 0.685 0.937 0.971 0.427 0.440 0.079 0.064 0.116 0.113 0.060 0.061
@@
TABLE V: VAE vs MVAE on MNIST Rotated across latent sizes. Best per cell is bold.
Latent MSE/test $\downarrow$ Probe Acc. $\uparrow$ Probe NLL $\downarrow$ Brier $\downarrow$ ECE $\downarrow$ NMI $\uparrow$ ARI $\uparrow$
VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE
2 0.030 0.030 0.668 0.672 1.000 0.989 0.451 0.448 0.029 0.030 0.268 0.215 0.175 0.134
4 0.029 0.029 0.759 0.757 0.755 0.746 0.339 0.336 0.037 0.031 0.235 0.254 0.142 0.156
16 0.046 0.045 0.116 0.116 2.295 2.295 0.898 0.898 0.015 0.012 0.006 0.005 0.002 0.002
32 0.029 0.029 0.711 0.714 0.883 0.879 0.397 0.394 0.025 0.027 0.245 0.231 0.154 0.134
256 0.040 0.040 0.271 0.264 2.083 2.090 0.822 0.823 0.061 0.057 0.082 0.082 0.047 0.049
512 0.031 0.030 0.757 0.749 0.754 0.780 0.339 0.349 0.028 0.028 0.240 0.243 0.154 0.159
@@
TABLE VI: VAE vs MVAE on MNIST Background Images across latent sizes. Best per cell is bold.
Latent MSE/test $\downarrow$ Probe Acc. $\uparrow$ Probe NLL $\downarrow$ Brier $\downarrow$ ECE $\downarrow$ NMI $\uparrow$ ARI $\uparrow$
VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE
2 0.081 0.081 0.789 0.801 0.649 0.615 0.301 0.286 0.029 0.031 0.436 0.466 0.316 0.340
4 0.081 0.081 0.851 0.853 0.480 0.464 0.222 0.218 0.038 0.035 0.426 0.479 0.303 0.354
16 0.083 0.083 0.545 0.546 1.247 1.234 0.571 0.567 0.075 0.068 0.443 0.430 0.262 0.248
32 0.081 0.081 0.816 0.820 0.576 0.568 0.266 0.261 0.036 0.033 0.433 0.456 0.314 0.346
256 0.081 0.081 0.712 0.703 0.855 0.885 0.397 0.410 0.043 0.044 0.464 0.427 0.311 0.268
512 0.082 0.081 0.851 0.857 0.475 0.454 0.219 0.211 0.031 0.029 0.446 0.470 0.299 0.343
@@
TABLE VII: VAE vs MVAE on MNIST Background Random across latent sizes. Best per cell is bold.
@@ V-B Quantitative Evaluation
Across rotated, background-images, and background-random settings, MVAE consistently strengthens calibration (NLL/Brier/ECE $\downarrow$) with comparable or better reconstruction (MSE $\downarrow$) and structure (NMI/ARI $\uparrow$). On MNIST Rotated (Table V), MVAE shows clear gains at mid latent sizes (e.g., z=4,16z{=}4,16) in probe accuracy and calibration, while very large $z$ yields small trade-offs (VAE edges on accuracy/NLL, MVAE on ECE). On MNIST Background Images (Table VI), results are near-ties with repeated MVAE calibration wins and negligible MSE deltas; atypical low-accuracy rows behave similarly for both models. On the noisier MNIST Background Random (Table VII), MVAE is consistently strong across $z$, improving probe accuracy, NLL/Brier, ECE, and often NMI/ARI.
@@
As shown in Table III, MVAE is competitive or better at nearly all $z$. Reconstruction is matched or slightly improved, while calibration metrics frequently favor MVAE (e.g., at z=4z{=}4 and z=16z{=}16 with NLL/Brier/ECE $\downarrow$). Linear-probe accuracy and clustering (NMI/ARI) exhibit modest but recurring MVAE gains, especially at mid (z=4z{=}41616) and large (z=512z{=}512) latent sizes.
@@
On CIFAR-10 (Table I), small $z$ shows mixed outcomes (MVAE improves calibration and structure; VAE may edge accuracy at z=2z{=}2). In the mid-$z$ regime (e.g., z=16z{=}16), MVAE improves accuracy and calibration concurrently; at very large $z$ (256–512), accuracy differences are marginal (often a slight VAE edge) while MVAE retains ECE/NLL/Brier advantages. On CIFAR-100 (Table II), MVAE yields small, steady gains across $z$ in accuracy and calibration, with NMI/ARI typically equal or slightly improved.

### v2

@@ Abstract
Learning latent representations that are simultaneously expressive, geometrically well-structured, and reliably calibrated remains a central challenge for Variational Autoencoders (VAEs). Standard VAEs typically assume a diagonal Gaussian posterior, which simplifies optimization but rules out correlated uncertainty and often yields entangled or redundant latent dimensions. We introduce the Multivariate Variational Autoencoder (MVAE), a tractable full-covariance extension of the VAE that augments the encoder with sample-specific diagonal scales and a global coupling matrix. This induces a multivariate Gaussian posterior of the form $\mathcal{N}(\mu_{\phi}(x),C\operatorname{diag}(\sigma_{\phi}^{2}(x))C^{\top})$, enabling correlated latent factors while preserving a closed-form KL divergence and a simple reparameterization path. Beyond likelihood, we propose a multi-criterion evaluation protocol that jointly assesses reconstruction quality (MSE, ELBO), downstream discrimination (linear probes), probabilistic calibration (NLL, Brier, ECE), and unsupervised structure (NMI, ARI). Across Larochelle-style MNIST variants, Fashion-MNIST, and CIFAR-10/100, MVAE consistently matches or outperforms diagonal-covariance VAEs of comparable capacity, with particularly notable gains in calibration and clustering metrics at both low and high latent dimensions. Qualitative analyses further show smoother, more semantically coherent latent traversals and sharper reconstructions. All code, dataset splits, and evaluation utilities are released to facilitate reproducible comparison and future extensions of multivariate posterior models.
@@ V-A Experimental Setup
Each configuration was defined by a dataset–model–latent-size triplet, ensuring a systematic exploration of both the baseline and proposed architectures. For every dataset, we trained models with latent dimensions ranging from low to high capacity (typically $d_{z}\in\{2,4,16,32,256\}$), enabling a consistent comparison across scales of representation.
Training was performed for $200$ epochs using the AdamW optimizer with a learning rate of $10^{-3}$ and a batch size of $100$ and an early stopping criteria. All experiments used fixed random seeds and stratified splits to ensure reproducibility. Evaluation was based on a comprehensive set of quantitative metrics: reconstruction quality via per-pixel mean squared error (MSE), generative quality via the evidence lower bound (ELBO), latent linear separability using a logistic regression probe, calibration performance through negative log-likelihood (NLL), Brier score, and expected calibration error (ECE), as well as unsupervised clustering measures using normalized mutual information (NMI) and adjusted Rand index (ARI).
Datasets include the Larochelle MNIST variants (basic, rotated, background images, background random), Fashion-MNIST, CIFAR-10, and CIFAR-100. For each dataset, both VAE and MVAE models are trained with latent sizes $\{2,4,16,32,256,512\}$ depending on input dimensionality.
@@
[…]
VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE
2 0.0352 0.0353 0.2185 0.2191 0.8708 0.8698 0.0422 0.0399 0.0744 0.0768 0.0355 0.0406
4 0.0294 0.0294 0.2696 0.2695 0.8409 0.8407 0.0392 0.0384 0.0937 0.0965 0.0500 0.0517
16 0.0186 0.0182 0.3392 0.3381 0.7947 0.7940 0.0392 0.0383 0.1126 0.1051 0.0668 0.0566
32 0.0154 0.0152 0.3644 0.3752 0.7705 0.7619 0.0293 0.0274 0.1112 0.1178 0.0582 0.0700
256 0.0149 0.0149 0.4188 0.4139 0.7217 0.7253 0.0335 0.0302 0.1076 0.1049 0.0567 0.0562
512 0.0153 0.0156 0.4379 0.4073 0.7047 0.7329 0.0344 0.0285 0.1027 0.1130 0.0537 0.0569
@@
TABLE I: VAE vs MVAE on CIFAR-10 across latent sizes. Bold indicates the better value per metric (↑ = higher is better, ↓ = lower is better).
[…]
VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE
2 0.0377 0.0376 0.0371 0.0364 0.9846 0.9843 0.0082 0.0067 0.1710 0.1729 0.0121 0.0122
4 0.0301 0.0298 0.0776 0.0769 0.9724 0.9735 0.0194 0.0199 0.2075 0.2070 0.0192 0.0190
16 0.0185 0.0183 0.1396 0.1408 0.9460 0.9458 0.0222 0.0229 0.2196 0.2163 0.0268 0.0256
32 0.0157 0.0154 0.1648 0.1678 0.9317 0.9304 0.0194 0.0214 0.2195 0.2201 0.0261 0.0277
256 0.0151 0.0152 0.1946 0.1958 0.9116 0.9116 0.0267 0.0288 0.2207 0.2239 0.0273 0.0268
512 0.0153 0.0153 0.2023 0.2009 0.9054 0.9084 0.0241 0.0246 0.2223 0.2230 0.0276 0.0265
@@
TABLE II: VAE vs MVAE on CIFAR-100 across latent sizes. Bold indicates the better value per metric (↑ = higher is better, ↓ = lower is better).
[…]
VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE
2 0.0270 0.0267 0.6523 0.6755 0.4824 0.4509 0.1003 0.0821 0.5501 0.5465 0.3758 0.3931
4 0.0193 0.0192 0.7421 0.7366 0.3589 0.3644 0.0275 0.0228 0.6162 0.5570 0.4425 0.4092
16 0.0144 0.0145 0.8024 0.7983 0.2765 0.2831 0.0173 0.0216 0.5069 0.5403 0.3538 0.3800
32 0.0144 0.0143 0.8209 0.8215 0.2545 0.2516 0.0238 0.0259 0.5278 0.5243 0.3842 0.3607
256 0.0151 0.0150 0.8587 0.8664 0.2056 0.1936 0.0301 0.0294 0.5390 0.5364 0.3834 0.3756
512 0.0157 0.0157 0.8649 0.8676 0.1967 0.1944 0.0277 0.0252 0.5299 0.5299 0.3776 0.3809
@@
TABLE III: VAE vs MVAE on Fashion MNIST across latent sizes. Bold indicates the better value per metric (↑ = higher is better, ↓ = lower is better).
[…]
VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE
2 0.0396 0.0395 0.4937 0.5470 0.6485 0.5657 0.1070 0.0732 0.4848 0.5177 0.3490 0.3701
4 0.0278 0.0278 0.8495 0.8490 0.2468 0.2499 0.0760 0.0786 0.6090 0.6060 0.5047 0.5082
16 0.0118 0.0117 0.8995 0.9094 0.1492 0.1363 0.0140 0.0115 0.5166 0.6459 0.3935 0.5902
32 0.0113 0.0111 0.9141 0.9219 0.1306 0.1159 0.0171 0.0165 0.6014 0.6367 0.5358 0.5668
256 0.0121 0.0120 0.9382 0.9520 0.0944 0.0751 0.0247 0.0202 0.5065 0.6166 0.3921 0.5342
512 0.0125 0.0128 0.9432 0.9486 0.0853 0.0777 0.0232 0.0211 0.5410 0.5256 0.4398 0.4062
@@
TABLE IV: VAE vs MVAE on MNIST Basic across latent sizes. Bold indicates the better value per metric (↑ = higher is better, ↓ = lower is better).
[…]
VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE
2 0.0483 0.0487 0.2025 0.2663 0.8686 0.7939 0.0809 0.0253 0.1739 0.1852 0.1345 0.0798
4 0.0371 0.0370 0.3406 0.3390 0.7251 0.7326 0.0122 0.0170 0.1631 0.1693 0.0835 0.0850
16 0.0159 0.0155 0.4934 0.4926 0.6284 0.6258 0.0246 0.0185 0.1201 0.1455 0.0643 0.0733
32 0.0130 0.0128 0.5291 0.5319 0.5952 0.5918 0.0214 0.0231 0.0914 0.1040 0.0479 0.0526
256 0.0136 0.0134 0.6748 0.6926 0.4532 0.4306 0.0806 0.0692 0.1281 0.0929 0.0661 0.0504
512 0.0143 0.0143 0.6925 0.6930 0.4358 0.4288 0.0822 0.0686 0.1285 0.1043 0.0655 0.0559
TABLE V: VAE vs MVAE on MNIST Rotated across latent sizes. Bold indicates the better value per metric (↑ = higher is better, ↓ = lower is better).
[…]
VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE
2 0.0451 0.0452 0.1164 0.1174 0.8983 0.8981 0.0138 0.0121 0.0054 0.0058 0.0017 0.0019
4 0.0401 0.0399 0.2650 0.2687 0.8213 0.8217 0.0530 0.0557 0.0851 0.0823 0.0493 0.0496
16 0.0300 0.0301 0.6634 0.6683 0.4540 0.4524 0.0301 0.0344 0.2571 0.2394 0.1679 0.1489
32 0.0287 0.0286 0.7106 0.7123 0.4002 0.3984 0.0267 0.0276 0.2559 0.2615 0.1634 0.1728
256 0.0296 0.0293 0.7535 0.7513 0.3455 0.3453 0.0340 0.0295 0.2483 0.2426 0.1579 0.1491
512 0.0308 0.0298 0.7548 0.7499 0.3428 0.3504 0.0284 0.0293 0.2427 0.2204 0.1585 0.1230
@@
TABLE VI: VAE vs MVAE on MNIST Background Images across latent sizes. Bold indicates the better value per metric (↑ = higher is better, ↓ = lower is better).
[…]
VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE VAE MVAE
2 0.0828 0.0828 0.5363 0.5465 0.5725 0.5682 0.0577 0.0647 0.4340 0.4429 0.2447 0.2654
4 0.0813 0.0811 0.7025 0.7033 0.4055 0.4050 0.0391 0.0444 0.4453 0.4334 0.2972 0.2800
16 0.0807 0.0808 0.7972 0.7985 0.2891 0.2854 0.0254 0.0276 0.4504 0.4683 0.3320 0.3583
32 0.0810 0.0806 0.8185 0.8153 0.2620 0.2665 0.0287 0.0318 0.4348 0.4557 0.3159 0.3283
256 0.0812 0.0810 0.8459 0.8552 0.2253 0.2101 0.0332 0.0311 0.4444 0.4618 0.3139 0.3316
512 0.0816 0.0815 0.8478 0.8535 0.2226 0.2151 0.0306 0.0310 0.4330 0.4597 0.3071 0.3341
@@
TABLE VII: VAE vs MVAE on MNIST Background Random across latent sizes. Bold indicates better value per metric (↑ higher is better, ↓ lower is better).
@@ V-B Quantitative Evaluation
Across the four Larochelle-style MNIST variants (basic, rotated, background images, and background random), the proposed MVAE generally matches or improves upon the baseline VAE, with the strongest gains appearing at low to moderate latent dimensions.
On MNIST Basic, MVAE achieves higher downstream accuracy at most latent sizes (e.g., $0.9520$ vs. $0.9382$ at z=256z{=}256 and $0.9486$ vs. $0.9432$ at z=512z{=}512) alongside lower Brier and ECE scores over a wide range of $d_{z}$ (e.g., Brier $0.0751$ vs. $0.0944$ and ECE $0.0202$ vs. $0.0247$ at z=256z{=}256). Correlated posteriors also sharpen cluster structure: at z=16z{=}16, NMI increases from $0.52$ to $0.65$ and ARI from $0.39$ to $0.59$; at z=32z{=}32, MVAE retains advantages in both NMI and ARI, whereas for very high $z$ (e.g., $512$) the VAE recovers a slight edge in clustering (NMI and ARI).
On MNIST Rotated, which introduces orientation noise, the picture is more mixed but still favorable overall. At z=16z{=}16 and $32$, MVAE reduces MSE (e.g., $0.0155$ vs. $0.0159$ at z=16z{=}16) and improves either accuracy or calibration: at z=32z{=}32, accuracy increases from $0.5291$ to $0.5319$ and Brier decreases from $0.5952$ to $0.5918$. ECE is substantially improved at z=2z{=}2 (from $0.0809$ to $0.0253$), while clustering metrics (NMI/ARI) alternate: MVAE is stronger at intermediate $z$ (e.g., NMI $0.1455$ vs. $0.1201$ at z=16z{=}16), whereas VAE retains higher NMI/ARI at large $z$ ($256$, $512$).
In MNIST Background Images, MVAE consistently improves or matches reconstruction and calibration across most latent sizes. For example, at z=32z{=}32 it achieves lower MSE ($0.0286$ vs. $0.0287$), lower Brier ($0.3984$ vs. $0.4002$), and comparable ECE, while also improving NMI and ARI ($0.2615$ vs. $0.2559$ and $0.1728$ vs. $0.1634$). At very high dimensions ($256$, $512$), MVAE continues to yield better MSE and ECE but VAE sometimes attains slightly higher clustering scores, suggesting that the benefits of covariance coupling are strongest at moderate capacities.
For MNIST Background Random, MVAE generally offers better reconstruction and classification at low and intermediate $z$ while trading off against calibration or clustering in some regimes. At z=2z{=}2, accuracy, Brier, NMI, and ARI all improve (e.g., accuracy $0.5465$ vs. $0.5363$, ARI $0.2654$ vs. $0.2447$), whereas at z=4z{=}4 the VAE retains an advantage in NMI/ARI despite slightly worse Brier. At higher $z$ (1616512512), both models perform similarly overall: MVAE often has lower Brier and higher NMI/ARI, while VAE can exhibit marginally better ECE or accuracy at specific latent sizes.
Overall, these results indicate that modeling full latent covariances via a coupling structure tends to enhance reconstruction, calibration, and clustering on the MNIST family, especially at low and moderate latent dimensionalities, while remaining competitive (and occasionally slightly worse) in certain high-dimensional clustering regimes.
@@
Fashion-MNIST, being visually richer than MNIST, exhibits more nuanced trade-offs between the two models. At very low latent size (z=2z{=}2), MVAE clearly outperforms VAE across most metrics: accuracy improves from $0.6523$ to $0.6755$, Brier from $0.4824$ to $0.4509$, ECE from $0.1003$ to $0.0821$, and ARI from $0.3758$ to $0.3931$ (with a marginal NMI decrease from $0.5501$ to $0.5465$). At z=16z{=}16, VAE achieves slightly better MSE and accuracy, whereas MVAE yields higher NMI and ARI ($0.5403$ vs. $0.5069$, $0.3800$ vs. $0.3538$), indicating more coherent clusters. For larger $z$ ($32$, $256$, $512$), the two models are largely comparable: MVAE often has slightly higher accuracy and lower calibration error (e.g., at z=512z{=}512, accuracy $0.8676$ vs. $0.8649$, Brier $0.1944$ vs. $0.1967$, ECE $0.0252$ vs. $0.0277$), while clustering metrics alternate between small benefits for VAE and MVAE. These outcomes suggest that the structured posterior primarily benefits calibration and clustering, with reconstruction and accuracy remaining close between models.
@@
On natural image datasets, MVAE demonstrates competitive or improved behavior, particularly in reconstruction and calibration, though gains are more modest than on binary domains.
For CIFAR-10, MVAE achieves lower or comparable MSE at intermediate latent dimensions (e.g., $0.0182$ vs. $0.0186$ at z=16z{=}16 and $0.0152$ vs. $0.0154$ at z=32z{=}32). Probe accuracy modestly improves at small and mid-range $z$ (e.g., $0.2191$ vs. $0.2185$ at z=2z{=}2 and $0.3752$ vs. $0.3644$ at z=32z{=}32), while VAE slightly outperforms MVAE at very high $z$ ($256$, $512$). Calibration metrics show more consistent gains: ECE is lower for MVAE at most sizes (e.g., $0.0274$ vs. $0.0293$ at z=32z{=}32), and Brier is marginally reduced. Clustering metrics (NMI/ARI) generally favor MVAE at low and intermediate $z$ (e.g., NMI $0.0768$ vs. $0.0744$, ARI $0.0406$ vs. $0.0355$ at z=2z{=}2; NMI $0.1178$ vs. $0.1112$, ARI $0.0700$ vs. $0.0582$ at z=32z{=}32), although VAE can be slightly better in a few configurations (e.g., NMI at z=16z{=}16).
For CIFAR-100, the task is substantially harder, and improvements are correspondingly smaller and more metric-dependent. MVAE consistently reduces MSE at most latent sizes (e.g., $0.0183$ vs. $0.0185$ at z=16z{=}16, $0.0154$ vs. $0.0157$ at z=32z{=}32), and achieves comparable or slightly better calibration on certain settings (e.g., Brier $0.9458$ vs. $0.9460$ at z=16z{=}16). Classification accuracy gains are modest and not strictly monotone (e.g., $0.1408$ vs. $0.1396$ at z=16z{=}16, $0.1678$ vs. $0.1648$ at z=32z{=}32, but slightly lower at z=2,4,512z{=}2,4,512). Clustering metrics again show small but interpretable differences: MVAE tends to match or slightly improve ARI at mid-range $z$ (e.g., $0.0277$ vs. $0.0261$ at z=32z{=}32), while VAE remains competitive or slightly better in NMI at some high-capacity settings.
Taken together, the CIFAR-10/100 results indicate that the MVAE’s gains persist outside grayscale domains, particularly for calibration and some aspects of cluster structure, while classification accuracy remains broadly similar between the two models and depends more strongly on latent size and dataset complexity.